# The life cycle of giant molecular clouds in simulated Milky Way-mass galaxies

### v1

@@ Abstract
Methods. We identify GMCs in simulated galaxies and track their evolution using cloud evolution trees. Via cloud evolution trees, we quantify the lifetimes and SFE of GMCs. We further apply our diagnostics on a suite of simulations with varying star formation and stellar feedback subgrid models and explore their impact together with galactic environments to the GMC life cycles.
Results. Our analysis reveals that GMCs undergo dynamic evolution, characterized by continuous gas accretion, gravitational collapse, and star formation, followed by disruption due to stellar feedback. The accretion process sustains the gas content throughout most of the GMC life cycles, resulting in a positive correlation between GMC lifetimes and their maximum masses. The GMC lifetimes range from a few to several tens of Myr, with two distinct dynamical modes: (1) GMCs near the galactic center experience strong tidal disturbances, prolonging their lifetimes when they remain marginally unbound; (2) those in the outer regions are less affected by tides, remain gravitationally bound, and evolve more rapidly. In all model variations, we observe that GMC-scale SFE correlates with the baryonic surface density of GMCs, consistent with previous studies of isolated GMCs. Additionally, we emphasize the critical role of galactic shear in regulating GMC-scale star formation and refine the correlation between local SFE and surface density by including its effects. These findings demonstrate how stellar feedback and galactic-scale dynamics jointly shape GMC-scale star formation in realistic galactic environments.
@@ 1 Introduction
Star formation is known to be inefficient on galactic scale, with a global molecular gas depletion time around $2-10\;\mathrm{Gyr}$, roughly one order of magnitude longer than the dynamical timescale of galactic disks and two orders of magnitude longer than the local free-fall time of giant molecular clouds (GMCs) (Kennicutt 1998; Bigiel et al. 2011; Kennicutt & Evans 2012; Sun et al. 2023). The long gas depletion timescale as well as low global star formation efficiency (SFE) in galaxies can be explained either by a uniformly low star formation rate across all individual GMCs (Krumholz & McKee 2005; Federrath & Klessen 2012; Pokhrel et al. 2021) or by a highly variable SFE during the evolution of GMCs with diverse GMC properties and galactic environments (Lee et al. 2016; Kruijssen et al. 2018; Chevance et al. 2023). To distinguish between the above two scenarios, it is critical to investigate the detailed life cycles of GMCs and the local star formation process observationally and theoretically.
@@
On the theoretical side, analytical models that consider the combined effects of gravitational collapse, cloud interaction, epicyclic motions, galactic shear, and large-scale gas streaming motions support the GMC lifetimes of several tens of Myr (Dobbs et al. 2015; Jeffreson & Kruijssen 2018). Numerical simulations of isolated GMCs also predict the lifetimes around a few free-fall times and reproduce the observed scatter in GMC-scale SFE, accounting for differences in GMC properties (Grudić et al. 2018; Li et al. 2019). However, both analytical models and idealized GMC simulations have notable limitations. For example, analytically predicted GMC lifetimes can be highly sensitive to assumptions about various evolutionary mechanisms, such as gravitational collapse, cloud-cloud collisions, or feedback-driven dispersal. These models often rely on simplified or isolated treatments of each mechanism, potentially overlooking complex interactions and dependencies between them. Furthermore, while numerical simulations capture several key physical processes within isolated GMCs, these studies (e.g. Murray et al. 2010; Dale et al. 2013; Skinner & Ostriker 2015; Howard et al. 2016; Kim et al. 2017; Grudić et al. 2018; Li et al. 2019; Grudić et al. 2021) ignore the critical effects on the galactic scale, such as shear from the larger galactic environment and continuous gas accretion from the surrounding medium. These environmental factors can significantly influence the stability, collapse timescale, and mass evolution of GMCs, highlighting the importance of including galactic context for more realistic predictions of GMC lifetimes and SFE variability.
@@
Over the past few decades, due to the advance in state-of-the-art galaxy formation simulations, especially the implementation of star formation (e.g. Cen & Ostriker 1992; Springel & Hernquist 2003; Li et al. 2017; Grudić et al. 2023; Girma & Teyssier 2024) and stellar feedback (e.g. Stinson et al. 2006; Vogelsberger et al. 2013; Ceverino et al. 2014; Hopkins et al. 2014; Smith et al. 2018; Marinacci et al. 2019) subgrid models, galaxy formation simulations start to be able to capture the structures of ISM and treat individual star-forming regions properly (Wang et al. 2015; Hopkins et al. 2018; Li et al. 2020, 2022; Reina-Campos et al. 2022; Nobels et al. 2023; Hopkins et al. 2023; Wibking & Krumholz 2023; Zhao et al. 2024). Some of these simulations include explicit stellar feedback from multiple channels—such as radiation, stellar winds, and supernovae—which, though not directly resolved, are crucial for modeling GMC properties and their evolution. As a result, these modern simulations now provide new laboratories to track the life cycles of GMCs (e.g. Benincasa et al. 2020; Jeffreson et al. 2021; Khullar et al. 2024) and investigate local SFE in realistic galactic environments.
@@
In this paper, we identify GMCs and track the temporal evolution during their life cycles in a set of simulations of isolated Milky Way-like galaxies introduced in Li et al. (2020, hereafter L20). This suite of simulations is performed with the moving-mesh code arepo (Springel 2010; Pakmor et al. 2016; Weinberger et al. 2020), using the Stars and MULtiphase Gas in GaLaxiEs (SMUGGLE) model (Marinacci et al. 2019, hereafter M19) to simulate key ISM processes that are key to the life cycles of GMCs. We investigate the lifetime distribution and SFE of the tracked GMC along time series, using the GMC identification in the three-dimensional space and establishing the novel tree network to follow the star formation within GMCs. The paper is organized as follows. In Section 2, we briefly introduce the physical ingredients in the simulations, describe the method to identify and track GMCs, demonstrate the construction of the cloud evolution tree, and illustrate the detection of star formation along the established tree network. In Section 3, we examine the properties of the GMCs, calculate their lifetime distribution, categorize the life cycle of the GMCs in different modes, and analyze the star formation process for different model variations. In Section 4, we compare our results with previous work and derive an analytical model for the SFE including both galactic scale effect and stellar feedback. Finally, in Section 5, we summarize the key results of this work.
@@ 2.1 Simulation details and model variations
The simulations analyzed in this work were introduced in L20 and are performed with the moving-mesh finite-volume hydrodynamic code arepo (Springel 2010; Pakmor et al. 2016; Weinberger et al. 2020). The simulations adopt the galaxy formation physics implemented in the SMUGGLE framework, a comprehensive galaxy formation model that includes radiative cooling and heating, star formation, and stellar feedback from radiation, stellar winds, and SNe (see M19). The initial conditions of the simulations are analog of a Milky Way-sized galaxy, which has a total mass of $\sim 1.6\times 10^{12}M_{\odot}$ and contains a stellar bulge and disc, a gaseous disc, and a dark matter halo. The total mass of the gaseous disk is approximately $\sim 9\times 10^{9}M_{\odot}$, with a mass resolution of $1.4\times 10^{3}M_{\odot}$ per gas cell. The stellar bulge has a total mass of $1.5\times 10^{10}M_{\odot}$ , resolved at $2.3\times 10^{3}M_{\odot}$ per stellar cell. Similarly, the stellar disk has a total mass of $4.73\times 10^{10}M_{\odot}$ with a mass resolution of $1.9\times 10^{3}M_{\odot}$ per stellar cell. We use adaptive gravitational softening for gas cells with a minimum softening length reaching $3.6$ pc, high enough to resolve the individual GMCs in the simulations. We also use a high-cadence snapshot write-out of 1 Myr to resolve and trace the temporal evolution of the clouds from birth to death.
The whole suite contains six different model variations with different combinations of stellar feedback channels and star formation efficiency, namely SFE1, SFE10, SFE100, Rad, SN, and Nofeed. The SFE1 is the fiducial run with $\epsilon_{\mathrm{ff}}=0.01$ and all stellar feedback channels, where $\epsilon_{\mathrm{ff}}$ is the star-formation efficiency per free-fall time, a factor that describes how fast a cold dense gas cell is converted into stars (see equation (7) in M19 for more details). The SFE10 and SFE100 runs are the same as SFE1, but with $\epsilon_{\mathrm{ff}}=0.1$ and $\epsilon_{\mathrm{ff}}=1$, respectively. The Nofeed, Rad, and SN runs have the same star formation parameters as SFE1, but with different feedback methods: Nofeed run turns off all stellar feedback channels; Rad run includes only radiative feedback; SN run includes only SN feedback. We refer the reader to L20 for a detailed description of the numerical setup of the simulations and the parameters of the subgrid physics.
@@ 2.2 Identifying GMCs
To capture the life cycle of GMCs in the simulated galaxies, we first identify individual GMCs in all snapshots for all six simulations. We use CloudPhinder [^1] https://github.com/mikegrudic/CloudPhinder, which starts from local density peaks in galaxies and then identifies the largest self-gravitating gaseous structures by searching for neighboring cells. We set the volumetric density threshold to $50\,\mathrm{cm}^{-3}$, similar to that of previous work (e.g. Tasker & Tan 2009; Grisdale et al. 2018; Fotopoulou et al. 2024), and only consider gas cells denser than this threshold to find candidate clumps. We set a fairly tolerant criterion of the virial parameter $\alpha\leq\alpha_{\mathrm{max}}=10$ as we aim to track clouds throughout their entire life cycles, where
$\alpha=\frac{2E_{\mathrm{kinetic}}}{|E_{\mathrm{potential}}|}=\dfrac{2\sum_{i}[m_{i}(|\@vec{v_{i}}-\@vec{v_{c}}|^{2}/2+c_{s,i}^{2})]}{\dfrac{1}{2}\sum_{i\neq j}\dfrac{Gm_{i}m_{j}}{|\@vec{r_{i}}-\@vec{r_{j}}|}},$ (1)
where $E_{\mathrm{kinetic}}$ and $E_{\mathrm{potential}}$ are the kinetic and potential energy of a candidate clump and ${v_{c}}$ is its center of mass velocity. $m_{i}$, $c_{s,i}$, ${v_{i}}$, and ${r_{i}}$ are the mass, sound speed, velocity, and position vector of the $i$-th gas cell within each GMC, respectively. To ensure that the identified GMCs are well resolved, we discard clouds that contain less than $10$ gas cells. To understand the physical properties in our GMC population and better compare the identified GMCs with observations and other theoretical work, we define effective radius $R_{\mathrm{eff}}$ and velocity dispersion $\sigma_{v}$ of GMCs as
$\begin{split}&R_{\mathrm{eff}}=\sqrt{\frac{5}{3}\frac{\sum(m_{i}r^{2}_{i})}{\sum m_{i}}},\\ &\sigma_{v}=\sqrt{\frac{\sum[m_{i}(|\@vec{v_{i}}-\@vec{v_{c}}|^{2}+c^{2}_{s,i})]}{\sum m_{i}}}.\end{split}$
@@ 2.3 Constructing GMC evolution network
Once we have identified GMCs across all epochs, we proceed to track their temporal evolution and construct an evolutionary network for each cloud. Similarly to Jeffreson et al. (2021, hereafter J21), we represent each identified cloud as a node that can connect with other nodes via edges. Unlike J21, which tracks GMCs using only the $x$ and $y$ coordinates of gas cells (essentially a two-dimensional approach), we use the full three-dimensional spatial information $(x,y,z)$ of the gas elements to determine connections between clouds identified in consecutive snapshots. To establish links between clouds across snapshots, we utilize the unique cell IDs of gas elements. Although the mesh reconstruction in arepo makes cell IDs not fully unique, many IDs remain consistent when the time interval between snapshots is short. This consistency enables the matching of gas elements across snapshots and helps identify connections between GMCs. Specifically, two clouds, $A$ and $B$, are linked if they share more than one common gas cell ID. This approach avoids assumptions about GMC movement trajectories required in a two-dimensional method and enables more accurate tracking in three dimensions. To exclude marginal connections, we apply an additional criterion: the total mass of shared gas elements, $M_{\mathrm{common}}$, must exceed $1$ percent of the mass of both clouds $A$ and $B$. Otherwise, the link between clouds $A$ and $B$ is discarded. After processing all clouds across snapshots, we produce a cloud evolution network consisting of nodes and edges for each simulation.
To further extract the lifespan of different clouds, we navigate through the network to construct cloud evolution trees, adhering to the following rules:
@@
As a result, each root node yields multiple potential trajectories originating from possible path choices within the multi-successor nodes. By multiplying the probabilities associated with each chosen path at each decision point, we can calculate the probability for all trajectories emerging from each root node. Specifically, for the tree traversing the node path $X_{0}\rightarrow X_{1}\rightarrow X_{2}\rightarrow...\rightarrow X_{n}$, we derive the probability as
$P(X_{0}\rightarrow X_{1}\rightarrow X_{2}\rightarrow...\rightarrow X_{n})=\prod_{0}^{n-1}P(X_{i}\rightarrow X_{i+1})$ (2)
Now we store the path probabilities for all possible evolution trees in our cloud network, based on which the star formation efficiency will be calculated in Section 2.4 and the distribution of the cloud lifetimes will be derived in Section 3.3. Note that our cloud evolution trees are different from those from the J21 approach in two key aspects:
@@
    The network in J21 considers any cloud with more descendants than progenitors as root nodes, which essentially treats all split clouds as new clouds. Here we start a new cloud evolution tree only when the cloud strictly has no progenitor. In this way, we can always capture the life cycles of GMCs from very early stages and follow their complete evolution.
@@
    The network in J21 gives equivalent path probabilities for all descendants in every multiple-descendant cloud, regardless of how much gaseous mass flows towards each descendant. Here we provide different weights to different descendants according to how much gaseous mass is flowing from the progenitor to each descendant. Thus, we focus on the majority of gaseous mass flow throughout cloud evolution trees.

### v2

@@ Abstract
Methods. We identified GMCs in simulated galaxies and tracked their evolution using cloud evolution trees. Via cloud evolution trees, we quantified the lifetimes and SFE of GMCs. We further applied our diagnostics to a suite of simulations with varying star formation and stellar feedback subgrid models and explored their impact together with galactic environments to the GMC life cycles.
Results. Our analysis reveals that GMCs undergo dynamic evolution, characterized by continuous gas accretion, gravitational collapse, and star formation, followed by disruption due to stellar feedback. The accretion process sustains the gas content throughout most of the GMC life cycles, resulting in a positive correlation between GMC lifetimes and their maximum masses. The GMC lifetimes range from a few to several tens of million years, with two distinct dynamical modes: (1) GMCs near the galactic center experience strong tidal disturbances, prolonging their lifetimes when they remain marginally unbound; (2) those in the outer regions are less affected by tides, remain gravitationally bound, and evolve more rapidly. In all model variations, we observe that GMC-scale SFE correlates with the baryonic surface density of GMCs, consistent with previous studies of isolated GMCs. Additionally, we emphasize the critical role of galactic shear in regulating GMC-scale star formation and refine the correlation between local SFE and surface density by including its effects. These findings demonstrate how stellar feedback and galactic-scale dynamics jointly shape GMC-scale star formation in realistic galactic environments.
@@ 1 Introduction
Star formation is known to be inefficient on a galactic scale, with a global molecular gas depletion time around $2-10\,\mathrm{Gyr}$, roughly one order of magnitude longer than the dynamical timescale of galactic disks and two orders of magnitude longer than the local free-fall time of giant molecular clouds (GMCs) (Kennicutt 1998; Bigiel et al. 2011; Kennicutt & Evans 2012; Sun et al. 2023). The long gas depletion timescale as well as low global star formation efficiency (SFE) in galaxies can be explained either by a uniformly low star formation rate across all individual GMCs (Krumholz & McKee 2005; Federrath & Klessen 2012; Pokhrel et al. 2021) or by a highly variable SFE during the evolution of GMCs with diverse GMC properties and galactic environments (Lee et al. 2016; Kruijssen et al. 2018; Chevance et al. 2023). To distinguish between the above two scenarios, it is critical to investigate the detailed life cycles of GMCs and the local star formation process observationally and theoretically.
@@
On the theoretical side, analytical models that consider the combined effects of gravitational collapse, cloud interaction, epicyclic motions, galactic shear, and large-scale gas streaming motions support the GMC lifetimes of several tens of million years (Dobbs et al. 2015; Jeffreson & Kruijssen 2018). Numerical simulations of isolated GMCs also predict the lifetimes around a few free-fall times and reproduce the observed scatter in GMC-scale SFE, accounting for differences in GMC properties (Grudić et al. 2018; Li et al. 2019). However, both analytical models and idealized GMC simulations have notable limitations. For example, analytically predicted GMC lifetimes can be highly sensitive to assumptions about various evolutionary mechanisms, such as gravitational collapse, cloud-cloud collisions, or feedback-driven dispersal. These models often rely on simplified or isolated treatments of each mechanism, potentially overlooking complex interactions and dependencies between them. Furthermore, while numerical simulations capture several key physical processes within isolated GMCs, these studies (e.g., Murray et al. 2010; Dale et al. 2013; Skinner & Ostriker 2015; Howard et al. 2016; Kim et al. 2017; Grudić et al. 2018; Li et al. 2019; Grudić et al. 2021) ignore the critical effects on the galactic scale, such as shear from the larger galactic environment and continuous gas accretion from the surrounding medium. These environmental factors can significantly influence the stability, collapse timescale, and mass evolution of GMCs, highlighting the importance of including galactic context for more realistic predictions of GMC lifetimes and SFE variability.
@@
Over the past few decades, due to the advance in state-of-the-art galaxy formation simulations, especially the implementation of star formation (e.g., Cen & Ostriker 1992; Springel & Hernquist 2003; Krumholz & Gnedin 2011; Padoan & Nordlund 2011; Semenov et al. 2016; Li et al. 2017; Dubois et al. 2021; Gensior et al. 2020; Valentini et al. 2023; Girma & Teyssier 2024; Ragone-Figueroa et al. 2024) and stellar feedback (e.g., Stinson et al. 2006; Vogelsberger et al. 2013; Ceverino et al. 2014; Hopkins et al. 2014; Smith et al. 2018; Lupi 2019; Marinacci et al. 2019; Keller et al. 2020) subgrid models, galaxy formation simulations start to be able to capture the structures of ISM and treat individual star-forming regions properly (Wang et al. 2015; Hopkins et al. 2018; Li et al. 2020, 2022; Reina-Campos et al. 2022; Nobels et al. 2023; Hopkins et al. 2023; Wibking & Krumholz 2023; Zhao et al. 2024). Some of these simulations include explicit stellar feedback from multiple channels—such as radiation, stellar winds, and supernovae—which, though not directly resolved, are crucial for modeling GMC properties and their evolution. As a result, these modern simulations now provide new laboratories to track the life cycles of GMCs (e.g., Benincasa et al. 2020; Jeffreson et al. 2021; Khullar et al. 2024) and investigate local SFE in realistic galactic environments.
@@
In this paper, we identified GMCs and tracked the temporal evolution during their life cycles in a set of simulations of isolated Milky Way-like galaxies introduced in Li et al. (2020, hereafter L20). This suite of simulations was performed with the moving-mesh code arepo (Springel 2010; Pakmor et al. 2016; Weinberger et al. 2020), using the Stars and MULtiphase Gas in GaLaxiEs (SMUGGLE) model (Marinacci et al. 2019, hereafter M19) to simulate key ISM processes that are crucial to the life cycles of GMCs. We investigated the lifetime distribution and SFE of the tracked GMC along time series, using the GMC identification in the three-dimensional space and establishing the novel tree network to follow the star formation within GMCs. The paper is organized as follows. In Section 2, we briefly introduce the physical ingredients in the simulations, describe the method to identify and track GMCs, demonstrate the construction of the cloud evolution tree, and illustrate the detection of star formation along the established tree network. In Section 3, we examine the properties of the GMCs, calculate their lifetime distribution, categorize the life cycle of the GMCs in different modes, and analyze the star formation process for different model variations. In Section 4, we compare our results with previous work and derive an analytical model for the SFE including both galactic scale effect and stellar feedback. Finally, in Section 5, we summarize the key results of this work.
@@ 2.1 Simulation details and model variations
The simulations analyzed in this work were introduced in L20 and were performed with the moving-mesh finite-volume hydrodynamic code arepo (Springel 2010; Pakmor et al. 2016; Weinberger et al. 2020). The simulations adopted the galaxy formation physics implemented in the SMUGGLE framework, a comprehensive galaxy formation model that includes radiative cooling and heating, star formation, and stellar feedback from radiation, stellar winds, and SNe. We refer the reader to M19 for more details of SMUGGLE. Here we highlight the sub-grid implementations that are relevant to this work. For star formation, star particles form from cold ($T<100\,\mathrm{K}$), dense ($n>100\,\mathrm{cm}^{-3}$), and self-gravitating molecular gas cells at a rate of:
$\dot{M}_{\star,\mathrm{cell}}=\epsilon_{\rm ff}\frac{M_{\mathrm{gas,cell}}}{t_{\mathrm{dyn,cell}}},$ (1)
where $\dot{M}_{\star,\mathrm{cell}}$, $\epsilon_{\mathrm{ff}}$, and $t_{\mathrm{dyn,cell}}=\sqrt{\frac{3\pi}{32G\rho_{\mathrm{cell}}}}$ are the star formation rate, SFE per free-fall time, and gas free-fall time at the cell scale, respectively. Here, $\rho_{\mathrm{cell}}$ is the density of a gas cell. The initial conditions of the simulations were analog of a Milky Way-sized galaxy, which had a total mass of $\sim 1.6\times 10^{12}M_{\odot}$ and contained a stellar bulge and disc, a gaseous disc, and a dark matter halo. The total mass of the gaseous disk was approximately $\sim 9\times 10^{9}M_{\odot}$, with a mass resolution of $1.4\times 10^{3}M_{\odot}$ per gas cell. The stellar bulge had a total mass of $1.5\times 10^{10}M_{\odot}$ , resolved at $2.3\times 10^{3}M_{\odot}$ per stellar cell. Similarly, the stellar disk had a total mass of $4.73\times 10^{10}M_{\odot}$ with a mass resolution of $1.9\times 10^{3}M_{\odot}$ per stellar cell. We used adaptive gravitational softening for gas cells with a minimum softening length reaching $3.6$ pc, high enough to resolve the individual GMCs in the simulations. We adopted high-cadence snapshot write-out for every $1$ Myr to trace the temporal evolution of the clouds. The cloud tracking algorithm is elaborated further in Section 2.3 and Section 2.4.
The whole suite contain six different model variations with different combinations of stellar feedback channels and star formation efficiency, namely SFE1, SFE10, SFE100, Rad, SN, and Nofeed. The SFE1 is the fiducial run with $\epsilon_{\mathrm{ff}}=0.01$ and all stellar feedback channels. The SFE10 and SFE100 runs are the same as SFE1, but with $\epsilon_{\mathrm{ff}}=0.1$ and $\epsilon_{\mathrm{ff}}=1$, respectively. The Nofeed, Rad, and SN runs have the same $\epsilon_{\mathrm{ff}}$ as SFE1, but with different feedback methods: Nofeed run turns off all stellar feedback channels; Rad run includes only early feedback [^1] We note that “early feedback” includes both the stellar winds and radiative feedback, and we use the label “Rad” for the early feedback-only run because the radiative feedback always dominates the early feedback.; SN run includes only SN feedback. We summarize the configuration of all model variations in Table 1 and refer the reader to L20 for a detailed description of the numerical setup of the simulations and the parameters of the subgrid physics.
Table 1: Summary of the six model variations analyzed in this work.
Name $\epsilon_{\mathrm{ff}}$ SN feedback Early feedback
SFE1 $0.01$ Yes Yes
SFE10 $0.1$ Yes Yes
SFE100 $1$ Yes Yes
Rad $0.01$ No Yes
SN $0.01$ Yes No
Nofeed $0.01$ No No
@@ 2.2 Identifying GMCs
To capture the life cycle of GMCs in the simulated galaxies, we first identified individual GMCs in all snapshots for all six simulations. We used CloudPhinder [^2] https://github.com/mikegrudic/CloudPhinder, which starts from local density peaks in galaxies and then identifies the largest self-gravitating gaseous structures by searching for neighboring cells. We set the volumetric density threshold to $50\,\mathrm{cm}^{-3}$, similar to that of previous work (e.g., Tasker & Tan 2009; Grisdale et al. 2018; Fotopoulou et al. 2024), and only considered gas cells denser than this threshold to find candidate clumps. We set a fairly tolerant criterion for the virial parameter $\alpha\leq\alpha_{\mathrm{max}}=10$ as we aimed to track clouds throughout their entire life cycles. We define the virial parameter of a given GMC as
$\alpha=\frac{2(E_{\mathrm{kin}}+E_{\mathrm{th}})}{|E_{\mathrm{grav}}|}.$ (2)
Here, $E_{\mathrm{kin}}$ and $E_{\mathrm{th}}$ are the kinetic and thermal energy of the cloud (measured in the cloud’s center-of-mass frame), and $E_{\mathrm{grav}}$ is the gravitational potential energy. The energies were computed as follows:
$E_{\mathrm{kin}}=\sum_{i}\frac{1}{2}m_{i}|\@vec{v_{i}}-\@vec{v_{c}}|^{2},$ (3)
$E_{\mathrm{th}}=\sum_{i}m_{i}u_{i},$ (4)
$E_{\mathrm{grav}}=-\frac{1}{2}\sum_{i\neq j}\frac{Gm_{i}m_{j}}{|\@vec{r_{i}}-\@vec{r_{j}}|},$ (5)
where $m_{i}$, $u_{i}$, ${v_{i}}$, and ${r_{i}}$ are the mass, internal energy per unit mass, velocity, and position of the $i$-th gas cell, respectively, and ${v_{c}}$ is the cloud’s center-of-mass velocity. $m_{j}$ and $\bm{r_{j}}$ are the mass and position of the $j$-th gas cell. To ensure that the identified GMCs are well resolved, we discarded clouds that contain fewer than ten gas cells [^3] We tested a threshold of up to $50$ cells and found that our main results—including cloud properties and lifetimes—remain robust. The only notable difference is the exclusion of some low-mass clouds at higher cell-number thresholds..
To understand the physical properties in our GMC population and better compare the identified GMCs with observations and other theoretical work, we define effective radius $R_{\mathrm{eff}}$ and velocity dispersion $\sigma_{v}$ of GMCs as
$\begin{split}&R_{\mathrm{eff}}=\sqrt{\frac{5}{3}\frac{\sum(m_{i}r^{2}_{i})}{\sum m_{i}}},\\ &\sigma_{v}=\sqrt{\frac{\sum[m_{i}(|\@vec{v_{i}}-\@vec{v_{c}}|^{2}+c^{2}_{s,i})]}{\sum m_{i}}},\end{split}$
where $c_{s,i}$ is the sound speed of of the $i$-th gas cell.
@@ 2.3 Constructing GMC evolution network
After we identified GMCs across all epochs, we proceeded to track their temporal evolution and constructed an evolutionary network for each cloud. Similarly to Jeffreson et al. (2021, hereafter J21), we represented each identified cloud as a “node” that could connect with other nodes via “edges”. We stress that each node represented a GMC identified at a specific snapshot. It was a bundle of information summarizing a GMC’s state at a given time. Edges were conceptual links between nodes across adjacent snapshots. Edges existed purely in the network topology and did not correspond to physical boundaries or geometrical features of the clouds. Unlike J21, which tracked GMCs using only the $x$ and $y$ coordinates of gas cells (essentially a two-dimensional approach), we used the full three-dimensional spatial information $(x,y,z)$ of the gas elements to determine connections between clouds identified in consecutive snapshots. To establish links between clouds across snapshots, we utilized the unique cell IDs of gas elements. Although a small fraction of cells might change their IDs across different epochs due to mesh refinement or de-refinement in arepo, we found that the majority retained consistent IDs when the time interval between snapshots was short, i.e., $1$ Myr. This consistency enabled the matching of gas elements across snapshots and helped identify connections between GMCs. Specifically, two clouds, $A$ and $B$, were linked if they shared more than one common gas cell ID. This approach avoided assumptions about GMC movement trajectories required in a two-dimensional method and enabled more accurate tracking in three dimensions. To exclude marginal connections, we applied an additional criterion: the total mass of shared gas elements, $M_{\mathrm{common}}$, must exceed one percent of the mass of both clouds $A$ and $B$. Otherwise, the link between clouds $A$ and $B$ was discarded. After processing all clouds across snapshots, we produced a cloud evolution network consisting of nodes and edges for each simulation.
To further extract the lifespan of different clouds, we navigated through the network to construct cloud evolution trees, adhering to the following rules:
@@
As a result, each root node yielded multiple potential trajectories originating from possible path choices within the multi-successor nodes. By multiplying the probabilities associated with each chosen path at each decision point, we calculated the probability for all trajectories emerging from each root node. Specifically, for the tree traversing the node path $X_{0}\rightarrow X_{1}\rightarrow X_{2}\rightarrow...\rightarrow X_{n}$, we derived the probability as
$P(X_{0}\rightarrow X_{1}\rightarrow X_{2}\rightarrow...\rightarrow X_{n})=\prod_{0}^{n-1}P(X_{i}\rightarrow X_{i+1}).$ (6)
Then, we stored the path probabilities for all possible evolution trees in our cloud network, based on which the SFE will be calculated in Section 2.4 and the distribution of the cloud lifetimes will be derived in Section 3.3. Note that our cloud evolution trees were different from those from the J21 approach in two key aspects:
@@
    The network in J21 considered any cloud with more descendants than progenitors as root nodes, which essentially treated all split clouds as new clouds. Here we started a new cloud evolution tree only when the cloud strictly had no progenitor. In this way, we could always capture the life cycles of GMCs from very early stages and follow their complete evolution.
@@
    The network in J21 gave equivalent path probabilities for all descendants in every multiple-descendant cloud, regardless of how much gaseous mass flowed toward each descendant. Here we provided different weights to different descendants according to how much gaseous mass was flowing from the progenitor to each descendant. Thus, we focused on the majority of gaseous mass flow throughout cloud evolution trees.